Humid air breaks down electrically when its molecules become ionized. This happens in an electric field of magnitude E = 3.0 ✕ 106 N/C. Calculate the magnitude of the electric force (in N) on an ion with a net single positive elementary charge in that field.
Answer #1

Here,

electric field , E = 3 *10^6 N/C

magnitude of electric force = E * charge

magnitude of electric force = 3 *10^6 * 1.602 *10^-19

magnitude of electric force = 4.81 *10^-13 N

the magnitude of electric force is 4.81 *10-13 N